A=1/2h(b1+b2) solve for b1
Answer by Boreal(15235)

multiply by 2
2a=h(b1+b2)
divide by h
(2a/h)=b1+b2
subtract b2
(2a/h)-b2=b1
There are other ways to write it, but this is acceptable